Current Yield for Annual Payments Heath Food Corporation’s bonds have 25 years remaining to maturity. The bonds have a face value of $1,000 and a yield to maturity of 7%. They pay interest annually and have a 7% coupon rate.

What is their current yield? Round your answer to two decimal places.
Answer #1

Information provided:

Face value= future value= $1,000

Time= 25 years

Yield to maturity= 7%

Coupon rate= 7%

Coupon payment= 0.07*1,000= $70

First, we need to compute the present value of the bond to arrive upon the current value.

The below has to be entered in a financial calculator to compute the present value of the bond:

FV= 1,000

N= 25

I/Y= 7

PMT= 70

Press the CPT key and PV to compute the present value.

The value obtained is 1,000.

Therefore, the current price of the bond is $1,000.

Current yield is calculated using the below formula:

Current Yield= Annual interest/Current price

                          = $70/ $1,000

                          = 0.07*100

                          = 7%     

Therefore, the current yield of the bond is 7%.

In case of any query, kindly comment on the solution